OpenStudy (anonymous):

Ball 1, with a mass of 20kg is moving to the right at 20 m/s. At what velocity should ball 2, with a mass of 40kg, move so that they both come to a standstill upon collision?

OpenStudy (rogue):

This question really belongs in the physics group, but here is how you do it... The balls will come to a standstill upon collision if they have the same momentum (going in opposite directions, so they cancel out). Momentum is given by the equation...\[P = mv\]Where P is momentum, m is mass, v is velocity. For your questions, since the momentum of ball 1 and ball 2 must have the same magnitude, you can set their momenta equal to each other.\[P_1 = P_2\]\[m_1 v_1 = m_2 v_2\]Substitute what you have and solve.\[20 kg * 20 m/s = 40 kg * v_2\]\[v_2 = \frac {400 kg*m/s}{40 kg} = 10 m/s\]So ball 2 should move at 10 m/s.